OpenStudy (anonymous):

I need help. Dropping a rock from a building that is 285 ft tall, how long would it take for it to hit the ground? if d=1/2gtE9 then what would the equation for time be?

OpenStudy (anonymous):

rad(2d/g)

OpenStudy (anonymous):

t= (2d/g)^1/2

OpenStudy (anonymous):

ok to be honest I'm in a physical science class with no calculus history. is that t=2d divided by g?

OpenStudy (anonymous):

all raised to the 1/2

OpenStudy (anonymous):

ok I don't think I did that right cause it's saying it would take 28.5 seconds to fall. If it falls at about 10m/s then it should take more like 9 1/2 seconds. I just don't know how to calculate it right

OpenStudy (anonymous):

are you sure that answer is correct? im getting about 4.2 secs. You dont want to use g = 9.81 m/s^2 here because your distance is measured in ft, not meters. you would want to use g = 32.174 ft/s^2 instead.

OpenStudy (anonymous):

\[285 = \frac{1}{2}g t^{2} \Rightarrow t = \sqrt{\frac{2*285}{32.174}}\]